# Zero-Shot Neural Architecture Search with Weighted Response Correlation

### v1

@@ 5 Discussion
Besides, following extensive literature reviews and experiments, we also have the same observation as the previous works [16, 30]. There is no single proxy that can consistently and robustly outperform other existing proxies in all situations. An optimal training-free proxy should be more comprehensive, considering various aspects of neural networks rather than just expressivity and generalizability. Furthermore, our experiments of voting proxies highlight the need for effective dynamic voting mechanisms, which is also an important direction for zero-shot NAS in the future.
@@
Although the zero-shot NAS algorithm we proposed has only demonstrated its effectiveness in image recognition, it has the potential to be extended to other downstream tasks in computer vision. Specifically, we can utilize the algorithm to design the optimal backbone architecture in image recognition and use the architecture in various downstream tasks. Moreover, our proposed proxy is agnostic to the architecture space and can be applied to other architectures (such as Transformers) or other fields (such as natural language processing and speech recognition).

### v2

@@ 5 Discussion
Besides, following extensive literature reviews and experiments, we also have the same observation as the previous works [19, 33]. There is no single proxy (including our proxies) that can consistently and robustly outperform other existing proxies in all situations. An optimal training-free proxy should be more comprehensive, considering various aspects of neural networks rather than just expressivity and generalizability. Furthermore, our experiments of voting proxies highlight the need for effective dynamic voting mechanisms, which is also an important direction for zero-shot NAS in the future.
@@
Although the zero-shot NAS algorithm we proposed has only demonstrated its effectiveness in image recognition, it has the potential to be extended to other downstream tasks in computer vision, and even other architectures and fields.